To estimate the size of the bear population on the Keweenan Peninsula, conversationalists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationalist's estimate of the size of the bear population?
Answer by stanbon(75887)

To estimate the size of the bear population on the Keweenan Peninsula, conversationalists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationalist's estimate of the size of the bear population?
---------------------------------
pop count/caught&tagged = caught including tagged/tagged
-------------------------------------------------------------
pop count/(50) = 100/2
pop count = 50(100/2) = 2500
==================================
Cheers,
Stan H.